# Can large language models generalize analogy solving like children can?

### v2

@@ 4 Discussion
This research received approval from the local ethics board on May 24, 2023 (Anonymized ID: XXX-XXX).
@@
The hypotheses, methods and analysis plan were preregistered for ethical approval. These were updated on July 26, 2023, prior to any data inspection or analysis of human data collection (conducted in June (children) and July (adults)), to accommodate new models and methods for LLM data collection which began Oct 12, 2023. There were minor deviations from the preregistration (for details see Appendix F).
@@
Materials, data, and analysis scripts for this manuscript are publicly available on the project’s Open Science Framework repository and linked GitHub code repository (Anonymized link: https://osf.io/jdty3/?view_only=8856260a39b74fcdb4f0204bbe909be1).
Appendix A Complete Base Item Set
Table 8 shows the complete base item set administered to all three participant groups.
Table 8:
Complete Base Item Set Administered to Adults, Children, and LLMs
Item ID Alphabet A B C D AB Rule AC Rule
A Practice a b j k successor(all, 1) shift
B Practice c d c d d j k j k k repetition(last, 2) shift
1 Latin a b a c g h g i successor(last, 1) shift
2 Latin c d c c e e m n m m o o successor(last, 1), repetition(all, 2) shift
3 Latin e f e h k l k n successor(last, 2) shift
4 Latin d e d f f g h g i i successor(last, 1), repetition(last, 2) shift
5 Latin c d b d m m n n l l n n predecessor(first, 1) shift, repetition(all, 2)
[…]
1 Symbol * @ * % ~ \$ ~ { successor(last, 1) shift
2 Symbol % ! % % ^ ^ = : = = ) ) successor(last, 1), repetition(all, 2) shift
3 Symbol @ % @ ^ # ~ # { successor(last, 2) shift
4 Symbol ! ^ ! # # \$ { \$ = = successor(last, 1), repetition(last, 2) shift
5 Symbol ^ # ! # = = : : { { : : predecessor(first, 1) shift, repetition(all, 2)
@@ Appendix B LLM prompt engineering results
As can be seen in Figure or Table 9, template 1, derived from Mitchell, (2021) worked best overall. Template 4, the best template found by Webb et al., (2023) worked well in Latin and Greek alphabets, but not as well for the Symbol list, which makes sense because $[[$ and ]]]] are symbols themselves. Our results are based on template 1.
Table 9:
Prompt Template Performance Mean Correct (SE) for Selected Models
@@ Appendix C LLM results without previous messages
As can be seen in Figure or Table 10, it was generally advantageous to include previous message history versus not. Of the LLMs we tested, there may be two possible exceptions to look out for in future work. Both Gemma-2 27B and Llama-3.1 405B had significantly higher accuracy (i.e., no overlapping SE margins) without message history on the Symbol alphabet. In both cases, the main result of lower performance on Greek and Symbol alphabets versus Latin alphabet still holds.
Table 10:
Performance Mean Correct (SE) for Selected Models by Including Message History vs Not Including Message History
@@
Appendix D Next-Previous Letter Task
The Next-Previous Letter Task was created to check whether LLMs were able to identify the previous and next two letters in an ordered sequence of letters. All items were administered one-by-one, without pre-pending previous conversation.
The LLM item instruction was:
    Here is an ordered list of letters or symbols
    ‘{Latin alphabet|Greek alphabet|Symbol list}’.
    Which letter or symbol is {one|two} {place|places} {before|after} {letter|symbol} ?
    Respond with only the letter or symbol.
The system instruction was:
    You are a helpful assistant that solves puzzles.
    Only give the answer, no other words or text.
Appendix E Ordered Symbol Task
We readministered the items from the template comparison (see Appendix B) to examine whether ordering the symbols by unicode value would improve the models’ performance on the Symbol alphabet. This did result in improved performance in Claude 3.5 and Gemma 2, where Claude 3.5 improved to the same performance level of the Greek alphabet. GPT-4o and Llama 3.1 did not improve from the reordering. In all cases our main finding - that performance degraded from Latin to the Greek and Symbol alphabets - still holds. However, in future experiments using a Symbol domain, it is important to realize that LLMs generally benefit by ordering symbols by unicode value.
Table 11: Means (SE) correct for each model on 5 items using 5 templates (25 items total). The last two columns represent performance using items where the Symbols were in the original order reported in the main analyses versus where the Symbol alphabet was ordered by unicode value.
@@ Appendix E Ordered Symbol Task
Appendix F Deviations from Pre-registration
In Table 12, we specify the differences between our OSF preregistration and the methods and analyses carried out in this study.
Table 12:
Differences Between OSF Preregistration and Present Study
OSF Preregistration Present Study
Models in LLM Sample
GPT-3 (text-davinci-003)
ChatGPT (GPT-4)
Claude-3.5
Gemma-2 27B
GPT-4o
Llama-3.1 405B
Scoring Method
Correct
Partially correct
Incorrect
Correct
Incorrect
Error categories in 3.5.2
LLM Instruction
Instructions included in the prompt
General and item instructions included in the prompt
System Instructions included in the system prompt

### v3

@@ Appendix A Open Science Practices
This study was approved by the ethics board at the University of Amsterdam, Social and Behavioural Sciences on May 24, 2023 (ID: FMG-2495).
@@
The hypotheses, methods and analyses were preregistered on the Open Science Framework (OSF) prior to ethical approval. These were updated on July 26, 2023, to accommodate new methods for LLM data collection. The main deviation from preregistration was using different LLMs than previously specified.
@@
Preregistration, materials, data, and code are publicly available on the project’s OSF repository: https://osf.io/jdty3/. A direct link to all data and code can be found here: https://github.com/cstevenson-uva/llm_letterstring_generalization/.
@@ Appendix B LLM prompt engineering results
As can be seen in Figure or Table 8, template 1, derived from Mitchell (2021) worked best overall. Template 4, the best template found by Webb et al. (2023) worked well in Latin and Greek alphabets, but not as well for the Symbol list, which makes sense because $[$ and ]] are symbols themselves. Our results are based on template 1.
Table 8: Prompt Template Performance Mean Correct (SE) for Selected Models
@@ Appendix C LLM results without previous messages
As can be seen in Table 9, it was generally advantageous to include previous message history versus not. Of the LLMs we tested, there may be two possible exceptions to look out for in future work. Both Gemma-2 27B and Llama-3.1 405B had significantly higher accuracy (i.e., no overlapping SE margins) without message history on the Symbol alphabet. In both cases, the main result of lower performance on Greek and Symbol alphabets versus Latin alphabet still holds.
Table 9: Mean Correct (SE) for LLMs with versus without Message History.
@@
Appendix D Ordered Symbol Task
We re-administered the items from the template comparison (see Appendix B) to examine whether ordering the symbols by unicode value would improve the models’ performance on the Symbol alphabet. This did result in improved performance in Claude 3.5 and Gemma 2, where Claude 3.5 improved to the same performance level of the Greek alphabet. For GPT-4o and Llama 3.1 there were no significant improvements from the reordering. In all cases our main finding —that performance degraded from Latin to the Greek and Symbol alphabets —still held. However, in future experiments using a Symbol domain, it is important to realize that LLMs generally benefit by ordering symbols by unicode value.
Table 10: Mean (SE) correct for LLMs with versus without Symbols were ordered by unicode value.